OpenStudy (vshiroky):

Solve the equation for y 2x+5y=10 y=-2/5x+2 Complete the ordered pair (5,__) (5,0)

jimthompson5910 (jim_thompson5910):

assuming you got \[\Large y = -\frac{2}{5}x+2\] for the first part you are correct for both parts

OpenStudy (vshiroky):

WOO!

OpenStudy (abmon98):

substitute the y into the first equation

OpenStudy (abmon98):

2x+5(-2/5x+2)=10, 2x-(10/5x+2)=10, 2x(5x+2)-10=10(5x+2), 10x^2-46+10 solve quadratic x=4.4, 0.22